# DustNET: enabling machine learning and AI models of dusty plasmas

### 23 Mar 2026 v2

@@ I Introduction
Dusty plasmas provide fertile ground to study and test multiscale (spatial and temporal), multi-phase plasma physics and plasma-material interactions, leveraging the fact that individual dust grains and their motion may be measured directly using imaging and other sensing methods Wang et al. (2020b). Dust absorption and emission of light in the form of interstellar reddening, which occur when dust particles absorb photons and re-emit photons at a longer wavelength, effectively place dust masks over stars and galaxies, and blur our vision of the universe and its evolution since the earliest time Li (2004); Khawaja et al. (2023).
@@
Central to the data-driven framework is to establish Dust Neural nEtworks Technology (DustNET), which, besides NN algorithms, also includes a curated dataset initiative modeled after ImageNet 166; J. Deng, W. Dong, R. Socher, L. Li, K. Li, and L. Fei-Fei (2009), designed to support the training and evaluation of DNNs for dusty plasmas. ImageNet, an open image database with more than 14 million labelled images, has been instrumental in advancing image classification, computer vision, deep learning, and AI in 2010s, a period which established the computational, algorithmic foundations of DNNs for ML and AI. Similar to ImageNet, images of dusty plasmas are available from almost every laboratory dusty plasma experiment, as described below. The existing experimental datasets, some of which have already been analyzed by neural networks, may form the initial base datasets of DustNET.
@@
To overcome these limitations, learned augmentation techniques have gained traction. For instance, Generative Adversarial Networks (GANs) Goodfellow et al. (2020) and Variational Autoencoders (VAEs) Diederik P Kingma (2013) can generate realistic synthetic data by learning the underlying distribution of the input data. These generative models are especially useful in domains like medical imaging or materials science, where acquiring labeled samples is costly. Recent advances in diffusion models have significantly raised the bar for generative quality and diversity. These models learn to reverse a noise process and generate high-fidelity samples conditioned on a prompt or class label Ho et al. (2020); Ramesh et al. (2022). Diffusion models such as Stable Diffusion, and transformer models such as DALL·E (see more below) are now widely used for image synthesis and have demonstrated strong performance in data-limited regimes. In scientific contexts, diffusion and transformer models are increasingly explored to generate physically plausible synthetic data to augment training sets Pinaya et al. (2022).
Another class of data-augmentation method is to use knowledge to supplement for large datasets. Physics knowledge is now combined with machine learning architecture in many different forms, for example, Physics-informed NN (PiNN) Raissi et al. (2019), Physics-guided NN (PgNN) Faroughi et al. (2023), physics-encoded neural networks (PeNNs), neural operators (NOs) Kovachki et al. (2023), and other physics-imbedding methods. Known physical laws or symmetry constraints are directly integrated into model training or data generation. Methods like Physics-Informed Neural Networks (PINNs) and simulation-informed data synthesis allow researchers to create synthetic datasets grounded in first principles. This ensures that augmented data respects the underlying behavior of physical systems, which is crucial for applications in materials science, fluid dynamics, and other domains where model interpretability and physical realism are essential. It has also been argued that the exceptional simplicity of physics-based functions hinges on properties such as symmetry, locality, compositionality and polynomial log-probability, and we explore how these properties translate into exceptionally simple neural networks approximating natural phenomena such as images Lin et al. (2017).
Recent years have seen rapid progress in generative artificial intelligence (GenAI) technology. Transformer-based architectures have revolutionized representation learning across modalities Vaswani et al. (2023). GenAI models, such as DALLE (from openAI), can generate realistic or imaginative images based on natural language descriptions, commonly known as ‘prompt engineering’. Other types of GenAI can create new content like text, images, audio, and video, based on the patterns and relationships it learns from vast amounts of training data. GenAI capability stems from significant advances in both model architectures such as diffusion models and transformers, and training strategies that involve billions or more model parameters. Some examples of using generative AI methods to generate images from text, will be given in examples below. In the near term, GenAI methods may be used for scientific visualization, education, and conceptual exploration. In the longer term, the GenAI methods may be used to augment simulation and experimental data, and also to validate NN models.
@@ II Diverse dusty plasma phenomena on different scales & Classification
Figure 1: Examples of dusty plasma dynamics on different temporal and spatial scales. On the smallest length scale ($L\leq r_{d}$), materials dynamics are important. The dust scale ($L\sim r_{d}$), based on direct measurements, can be used to examine dust motion using Newton’s laws. The largest scale ($L>r_{d}$) shows rich structures and its evolution, ranging from Coulomb crystals, dust waves to planetary rings, interstellar and intergalactic dust clouds. The thumbnail inserts are from the web searches or scientific literature, such as Boufendi and Bouchoule (1994); Winter (2004); Ticoş et al. (2021); Merlino (2021); Guo et al. (2022).
Naturally occurring dusty plasmas on Earth result from interactions between dust particles, originating from both natural sources and human activities, such as rocket launches, and atmospheric or ionospheric plasmas. Examples include noctilucent clouds (also known as polar mesospheric clouds), auroral plasmas that produce the northern and southern lights, and lightning storms. The noctilucent clouds, with a size of 10 to 100 km that scatter sunlight, has an estimated electron density ($n_{e}$) of 103 - 105 cm-3, a dust density $n_{d}\sim$ 10-3 - 10-1 cm-3 and the average dust diameter about 0.3 µm. Temperatures below 150 K or -120 oC are required to form noctilucent clouds. On the planetary and space scale, dusty plasmas, such as comet tails, Saturn’s rings and spokes Horányi et al. (2004); Merlino (2021), can now be observed at close encounters or even in-situ through lunar and planetary missions Wahlund et al. (2009). On the galactic scale, cosmic dusty plasmas are now routinely observed by using the Earth-bound and space telescopes Mendis and Rosenberg (1994). Lately, the James Webb Space Telescope has revealed intricate details of interstellar dust and gas clouds, allowing astronomers to map the three-dimensional structures of the interstellar medium for the first time 228.
@@ II.2 Newtonian dynamics
There are some nontrivial challenges in integrating an ‘optical tweezer’ with a dusty plasma experiment. All optical designs and configurations must be tailored to fit within the predefined space and geometry of the plasma system while ensuring sufficient optical trapping robustness and efficiency. Additionally, we must consider unavoidable mechanical vibrations and potential electromagnetic interference, which affect the optical trapping system. In our first optical trapping experiment in the MDPX system E. Thomas et al. (2012), as shown in Fig. 4(a), single particles were successfully trapped. Specifically, 6.2 µm glass spheres were introduced into the plasma and trapped near the geometric center of the plasma volume above the electrode surface, as shown in Fig. 4(b, c), in which (b) and (c) are the side and top views respectively. In addition to the optically trapped particle (circled), there are some particles trapped in the sheath, as viewed from the top. During this ten-day experiment, no attempts were made to manipulate trapped particles in the plasma, nor were measurements of plasma parameters carried out using the trapped particle. Future developments are focused on upgrading the optical configuration of the trapping system. Nevertheless, this work demonstrates the feasibility of integrating an optical trapping system into an existing plasma facility and successfully trapping single particles in the plasma.
@@
Electric field $E$ and particle charge $Q_{d}$ are two crucial parameters in dusty plasma, which influence key plasma properties such as electron, ion, and neutral atom or molecule densities, electron temperature, electric potential, and sheath structure. In many cases, $E$ at a location where a charged particle resides can be determined given a known $Q$, and vice versa. When a charged particle is optically trapped within a dusty plasma, its subsequent motions are primarily governed by electrostatic force $F_{e}$, gravitational force $F_{g}$, ion drag force $F_{id}$, and neutral drag force $F_{nd}$ when optical trapping forces are off. Assuming negligible speeds of the particle and the absence of other nearby particles, $F_{id}$ and $F_{nd}$ can be neglected. As $F_{g}$ remains unchanged since the particle’s mass is constant and $F_{e}$ varies proportionally with $E$ as $Q$ is a constant within a small region in the plasma, the particle’s motions are driven by changes in $E$. Following this principle, we trapped a single particle in a specific location and recorded its subsequent motions when the trapping force was off. Using a known particle charge $Q$ derived from a similar particle (same size and material) charged in a RF plasma under identical operational conditions (input RF power, plasma gas, pressure, and configuration), we measured $E$ along the particle’s trajectory. Fig.6(a) presents the recoded particle’s motion trajectories in the x-z plane, and Fig.6 (b) shows results resulting from the analysis of the x- and z- components of each trajectory, leading to corresponding $E_{x}$ and $E_{z}$ components. Two sets of data in Figs. 6(a) and (b) are from the repeated measurements involved in trapping a single particle in the approximately same (x, z) location where the particle started its motions after the trapping force was off. The data in (a) and (b) show good reproducibility. The measured E distributions have a spatial resolution of up to tens of microns along the z-axis. Correspondingly, the precision of E measurements exceeds 2 V/m. Note that the trapped-and-released particle travels in a longer distance along the x-axis than along the z-axis. This disparity is because $E_{x}$ remains approximately constant in the traveled range along the x-axis, while $E_{z}$ changes drastically in the z-direction, as illustrated in Fig. 6(b)Takizawa and Sasaki (2004); James et al. (2005).
Figure 6: High precision mapping electric field $E$ around a localized position using a trapped particle, with a space resolution of micrometers. (a) Recorded motion trajectories of the single trapped particle when the trapping force was off, (b) measured $E_{z}$ distribution along the z-axis. The inset in (b) shows that $E_{x}$ is nearly the same in the range of 5000 µm. The color and symbol denote repeated measurements.
@@ II.3 Structure dynamics
Figure 7: A picture taken during the last PK-3 Plus experiment on the International Space Station. Some “large” particles of about 1 mm in diameter were induced into the experiment chamber by accident. Interesting interactions between the large particles with the smaller particles were observed. Image credit: The Max Planck Institute for Extraterrestrial Physics (MPE).
@@ III Dusty plasma models
Computational plasma physics is also sometimes called ‘numerical or computational experiments’ Tajima (2004). Computational dusty plasma physics consists of plasma simulations, dust simulations, and dust-plasma interaction simulations. As computing power, memory and algorithms continue to improve, simulations of a full plasma experiment as large as the ITER scale with high fidelities, which usually combines multiple codes to resolve and connect phenomena across scales, is now a reality. WDMApp 430, for example, couples two advanced and highly scalable gyrokinetic codes, XGC and GENE (or GEM), where the former is a particle-in-cell code optimized for the treating the edge plasma while GENE (GEM) is a continuum code optimized for the core. In another example, Fusion Energy Reactor Models Integrator (FERMI) couples plasma (MHD) simulations with structural mechanics (first wall, blanket designs), neutronics, and fluid flows Badalassi et al. (2023). In practice, many ‘phenomenal approaches’ or reduced order models are adopted, which require isolation of important physics, simplified numerical algorithms, aiming for numerical results that can be directly compared with measurements or observations. Neural networks offer a completely new way to develop pragmatic reduced order models.
@@ III.1 Plasma models
As a branch of physics, the basic multi-physics framework for plasmas, which includes Newton’s mechanics, classical mechanics, Maxwell’s theory of electromagnetism, relativity, thermodynamics, statistical physics and kinetic theory, had been established by the 1950s for most practical purposes, as summarized in Fig. 8. Quantum mechanics, which has often been neglected in plasma physics, is growingly incorporated into the existing ‘classical plasma physics’ framework to better understand plasmas under extreme conditions or when the microscopic physics details of individual electrons, atoms or molecules are needed, such as in plasma chemistry. Many advances in plasma physics require modeling of very complex systems in laboratory or nature. Examples, as highlighted in Fig. 8, are magnetic fusion plasmas, space plasmas, laser plasmas, plasma accelerators, and laser fusion. One of the latest accomplishments happened at the National Ignition Facility (NIF), where nuclear fusion ignition was demonstrated in 2022. The ‘devils’ of plasma modeling have been hiding in the ‘computational details or high fidelities’ to this day. Whole device modeling of fusion plasmas, such as WDMApp, has recently become possible due to exascale computing 430. Advances in machine learning and neural networks in the 2010s, mostly due to advances outside plasma physics, is a growingly adopted in plasma physics to further enhance the traditional high-performance computing.
@@
In the context of dusty plasmas, one may divide plasma models into reactive and non-reactive plasma models. The main difference between a reactive and a non-reactive plasma lies in the significance of plasma chemistry Kushner (1988); Hollenstein (2000); Fridman (2008). In a non-reactive dusty plasma, plasma chemistry may be ignored, and dust mass, geometry, and chemical composition stay constant Pustylnik et al. (2021). We may call such models as ‘constant dust’ models. Modeling plasma chemistry can capture the material dynamics such as dust growth and destruction, meanwhile, it requires a large number of chemical species, radicals, chemical reaction cross sections and reaction rates. In industrial plasma settings Fridman (2008), including high-temperature plasma boundaries, modeling plasma chemistry is almost always necessary. Open source software, such as OpenFOAM 298, together with commercial software, such as VSimPlasma 410 and COMSOL 167, is available for the complicated physics and chemistry, and for integrated multi-disciplinary system modeling and optimization Martins and Ning (2021). In many laboratory low-temperature plasmas, non-reactive plasma assumption may be adopted for simplified computation of Newtonian dynamics and structure dynamics, while ignoring materials dynamics and plasma chemistry. For each kind of particle (electron, ions, atoms, dust grains), atomistic, kinetic, or fluid models may be used, depending on the number of particles ($N$) involved and the collisional properties of the particle with the rest of system, as shown in Fig. 9. At a more fundamental level, the atomistic or first-principle models simulate individual particle motion and their interactions with others, such as particle-particle interactions, particle - electromagnetic field interactions, and particle-wall interactions. The computational cost for the best atomistic codes scales polynomially with $N$ as $O(N)$, which may not be practical when $N>10^{12}$ assuming Newtonian dynamics (also other parameters, such as time-step size, duration of the simulation, memory, and communications) on the state-of-the-art supercomputers, such as the Frontier supercomputer in the Oak Ridge National Laboratory 34.
@@
Besides computational cost, the selection of atomistic, kinetic or statistical, or fluid or continuum model for a specific type of particle also depends collisional properties of the particle. The collisional properties may be measured by the Knudsen number $Kn$, the ratio of the collisional mean free path $\lambda$ to the characteristic length $a$. KnKn for ions interaction with a dust particle with a radius $r_{d}$ may be defined as $Kn_{d}=\lambda/r_{d}$, while $Kn_{p}=\lambda/L$ for the plasma as whole, with $L$ being the plasma size and $r_{d}\ll L$ is very common in laboratory and nature. In other words, while ions may be modeled as fluid when the problem is to quantitatively calculate the global plasma properties such as transport properties or velocity field, ion-dust interaction may need kinetic models if $Kn_{d}>1$.
@@
Attempts at balancing pragmatic computationally accessible model for a dusty plasma with scientific and logical rigor will necessarily run into the famous Hilbert’s sixth problem Slemrod (2015); Gorban (2018); Deng et al. (2025), which aimed at establishing a rigorous mathematical framework for all the laws of physics (which, of course, includes dusty plasma physics and computation), and axiomatizing the laws of physics like what has been done to geometry. From atomistic models to kinetic models, such as the Boltzmann equation, for example, the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) equations hierarchy reduces the comprehensive $N$-particle description in phase space to the single-particle distribution as in the Boltzmann equation Harris (1971). Another reduction is from the kinetic equation to the fluid equations, as in Fig. 9, through the Chapman-Enskog method for charge-neutral gas. Equivalent approaches to the Chapman-Enskog method, often known as solving the closure problem, remain an active field of research in and outside plasma physics Kogelbauer and Karlin (2025).
@@
As examples of ‘atomistic’ or ‘$N$-particle’ models, particle-in-cell (PIC) in combination with Monte Carlo collision (MCC), or ‘PIC-MCC’ method, is widely used in simulations of partially ionized gases Birdsall (1991); Vahedi and Surendra (1995); Alves et al. (2018). The number of particles $N$ that can be simulated continues to increase from 103 in the 1970s to above 106 Birdsall and Langdon (2004); Verboncoeur (2005). In PIC methods, real particles (like electrons and ions) are represented by computational particles, also called macro-particles or super-particles, each representing a large number of real particles (typically 105–108Hara (2019). A weight factor is used to relate the density of computational particles to the physical density of the electrons or ions Hockney and Eastwood (1988). Axisymmetric cylindrical or 2D PIC simulations have been applied to complex boundary conditions that match those found in an experimental discharge chamber, such as a GEC cell, or to DC discharge tubes, such as the PK-4 experiment onboard the International Space Station Donkó (2011); Hartmann et al. (2020). The MCC technique is based on elementary collisions of electrons, ions and fast neutral atoms with the thermal background gas and is able to describe charged particle and energy transport at the kinetic level for individual particles. The PIC-MCC method provides a fully self-consistent, kinetic level description of the gas discharge along with space and time resolved properties (including electric potential, densities, fluxes, energy distributions, surface processes) across the entire discharge including the walls of the discharge. A hybrid-PIC-MCC (h-PIC-MCC) algorithm was described to study the effect of dust charge fluctuation on ion acoustic waves in a dusty plasma Changmai and Bora (2020).
@@
Molecular dynamics (MD) are often used to simulate $N$-atom/molecule/ion/dust (microparticle) dynamics. The key idea behind a MD simulation is straightforward by using Newton’s laws of motion to trace particle motion Neyts and Brault (2017); Hollingsworth and Dror (2018). While PIC simulations track computational particles and their interactions with fields, MD simulations treat directly real particles and their interactions, including interatomic potentials such as the Lennard-Jones potential, Coulomb forces, or bonded force models. MD simulations do not require a spatial grid for force calculations unless coupled with external fields; forces are computed directly from the particle positions Frenkel and Smit (2023). One approach is to adopt existing MD codes for dusty plasmas, e.g. the microparticles were modeled using the classical MD package Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Schwabe and Graves (2013). Another approach is to develop customized MD codes for specific dusty plasma problems, e.g, molecular dynamics simulation of ion flows around microparticles was implemented in a code called molecular asymmetric dynamics (MAD) and its variants Piel (2017a); Piel et al. (2018a, b). In another example, Dynamic Response of Ions and Dust (DRIAD) is a MD code for ions and dust Matthews et al. (2020), where the dust and ion motion are resolved at different time steps, with $\Delta t=$ 10-4 s for dust and 10-9 s for ions.

### 09 Apr 2026 v3

@@ I Introduction
Dusty plasmas provide fertile ground to study and test multiscale (spatial and temporal), multi-phase plasma physics and plasma-material interactions, leveraging the fact that individual dust grains and their motion may be measured directly using imaging and other sensing methods Wang et al. (2020). Dust absorption and emission of light in the form of interstellar reddening, which occur when dust particles absorb photons and re-emit photons at a longer wavelength, effectively place dust masks over stars and galaxies, and blur our vision of the universe and its evolution since the earliest time Li (2004); Khawaja et al. (2023).
@@
Central to the data-driven framework is to establish Dust Neural nEtworks Technology (DustNET), which, besides NN algorithms, also includes a curated dataset initiative modeled after ImageNet 176; J. Deng, W. Dong, R. Socher, L. Li, K. Li, and L. Fei-Fei (2009), designed to support the training and evaluation of DNNs for dusty plasmas. ImageNet, an open image database with more than 14 million labelled images, has been instrumental in advancing image classification, computer vision, deep learning, and AI in 2010s, a period which established the computational, algorithmic foundations of DNNs for ML and AI. Similar to ImageNet, images of dusty plasmas are available from almost every laboratory dusty plasma experiment, as described below. The existing experimental datasets, some of which have already been analyzed by neural networks, may form the initial base datasets of DustNET.
@@
[…]
Another class of data augmentation approaches incorporates domain knowledge, including physics, geometry, topology, and statistical principles, to supplement and compensate for limited labelled data availability. Physics knowledge is now combined with machine learning architecture in many different forms, for example, Physics-informed NN (PiNN, PINN) Raissi et al. (2019), Physics-guided NN (PgNN) Faroughi et al. (2023), physics-encoded neural networks (PeNNs), neural operators (NOs) Kovachki et al. (2023), and other physics-imbedding methods. Known physical laws or symmetry constraints are directly integrated into model training or data generation. Methods like PINNs and simulation-informed data synthesis allow researchers to create synthetic datasets grounded in first principles. This ensures that augmented data respects the underlying behavior of physical systems, which is crucial for applications in materials science, fluid dynamics, and other domains where model interpretability and physical realism are essential. It has also been suggested that the exceptional simplicity of physics-based functions hinges on properties such as symmetry, locality, compositionality and polynomial log-probability, and these properties can be translate into exceptionally simple neural networks approximating images of natural phenomena Lin et al. (2017).
[…]
@@ II Diverse dusty plasma phenomena on different scales & Classification
Figure 1: Examples of dusty plasma dynamics on different temporal and spatial scales. On the smallest length scale ($L\leq r_{d}$), materials dynamics are important. The dust scale ($L\sim r_{d}$), based on direct measurements, can be used to examine dust motion using Newton’s laws. External forces using lasers, electron beams (e-beam), and plasma flows have been reported. The largest scale ($L>r_{d}$) shows rich structures and their evolution, ranging from Coulomb crystals, dust waves to planetary rings, interstellar and intergalactic dust clouds. The thumbnail inserts are from the web searches or scientific literature, such as Boufendi and Bouchoule (1994); Winter (2004); Ticoş et al. (2021); Merlino (2021); Guo et al. (2022).
Naturally occurring dusty plasmas on Earth result from interactions between dust particles, originating from both natural sources and human activities, such as rocket launches, and atmospheric or ionospheric plasmas. Examples include noctilucent clouds (also known as polar mesospheric clouds), auroral plasmas that produce the northern and southern lights, and lightning storms. The noctilucent clouds, with a size of 10 to 100 km that scatter sunlight, has an estimated electron density ($n_{e}$) of 103 - 105 cm-3, a dust density $n_{d}\sim$ 10-3 - 10-1 cm-3 and the average dust diameter about 0.3 µm. Temperatures below 150 K or -120 oC are required to form noctilucent clouds. On the planetary and space scale, dusty plasmas, such as comet tails, Saturn’s rings and spokes Horányi et al. (2004); Merlino (2021), can now be observed at close encounters or even in-situ through lunar and planetary missions Wahlund et al. (2009). On the galactic scale, cosmic dusty plasmas are now routinely observed by using the Earth-bound and space telescopes Mendis and Rosenberg (1994). Lately, the James Webb Space Telescope has revealed intricate details of interstellar dust and gas clouds, allowing astronomers to map the three-dimensional structures of the interstellar medium for the first time 239.
@@ II.2 Newtonian dynamics
There are some nontrivial challenges in integrating an ‘optical trapping’ with a dusty plasma experiment. All optical designs and configurations must be tailored to fit within the predefined space and geometry of the plasma system while ensuring sufficient optical trapping robustness and efficiency. Additionally, we must consider unavoidable mechanical vibrations and potential electromagnetic interference, which affect the optical trapping system. In our first optical trapping experiment in the MDPX system E. Thomas et al. (2012), as shown in Fig. 4(a), single particles were successfully trapped. Specifically, 6.2 µm glass spheres were introduced into the plasma and trapped near the geometric center of the plasma volume above the electrode surface, as shown in Fig. 4(b, c), in which (b) and (c) are the side and top views respectively. In addition to the optically trapped particle (circled), there are some particles trapped in the sheath, as viewed from the top. During this ten-day experiment, no attempts were made to manipulate trapped particles in the plasma, nor were measurements of plasma parameters carried out using the trapped particle. Future developments are focused on upgrading the optical configuration of the trapping system. Nevertheless, this work demonstrates the feasibility of integrating an optical trapping system into an existing plasma facility and successfully trapping single particles in the plasma.
@@
Electric field $E$ and particle charge $Q_{d}$ are two crucial parameters in dusty plasma, which influence key plasma properties such as electron, ion, and neutral atom or molecule densities, electron temperature, electric potential, and sheath structure. In many cases, $E$ at a location where a charged particle resides can be determined given a known $Q$, and vice versa. When a charged particle is optically trapped within a dusty plasma, its subsequent motion is primarily governed by the electrostatic force $F_{e}$, gravitational force $F_{g}$, ion drag force $F_{id}$, and neutral drag force $F_{nd}$ when the optical trapping forces are turned off. Over the small spatial region explored by the particle, its charge was assumed to remain approximately constant. Using a charge value obtained from an identical particle under the same RF plasma conditions, we reconstructed the electric field along the particle’s trajectory Ekanayaka et al. (2025). Because the sheath electric field is strongly axial, the axial electrostatic force nearly balances gravity. As a result, the particle exhibits minimal axial displacement, and the transverse direction dominates its motion, as shown in Fig. 6(a). After release, the particle travels with nearly constant axial velocity, enabling determination of the axial electric field from the balance of gravitational, electrostatic, and ion drag forces. Turning off the optical trapping forces is assumed not to perturb the particle charge. Although the particle undergoes millimeter-scale transverse motion, previous studies indicate that dust charge depends mainly on axial position and is only weakly sensitive to transverse variations Schleede et al. (2018). In our measurements, axial displacement is limited to only a few micrometers, making charge variation negligible. The resulting axial electric field is on the order of $10^{3}\,\mathrm{V/m}$, consistent with typical sheath values. Figure 6(a) shows the particle trajectory in the xzxz plane, while Fig. 6(b) presents the corresponding transverse electric field $E_{x}$. Two repeated measurements with similar initial conditions $(x,z)$ demonstrate good reproducibility. The extracted electric-field distributions achieve spatial resolution on the order of tens of micrometers and precision better than $1\,\mathrm{V/m}$. The particle’s significantly larger displacement along the $x$-axis than along the $z$-axis confirms that $E_{x}$ is relatively small and nearly uniform, as illustrated in Fig. 6(b). In contrast, $E_{z}$ dominates the force balance, resulting in the observed minimal axial motion.
Figure 6: High precision mapping of the electric field E around a localized position using a trapped particle, with a space resolution of micrometers. (a) Recorded motion trajectories of the single trapped particle when the trapping force was off, with color indicating time progression, (b) measured Ex distribution along the x-axis. The color and symbol denote repeated measurements.
@@ II.3 Structure dynamics
Figure 7: A picture taken during the last PK-3 Plus experiment on the International Space Station. Some “large” particles of about 1 mm in diameter were induced into the experiment chamber by accident. Interesting interactions between the large particles with the smaller particles were observed. Image credit: The Max Planck Institute for Extraterrestrial Physics (MPE). Additional information may be found in Schwabe et al. (2017).
@@ III Dusty plasma models
Computational plasma physics is also sometimes called ‘numerical or computational experiments’ Tajima (2004). Computational dusty plasma physics consists of plasma simulations, dust simulations, and dust-plasma interaction simulations. As computing power, memory and algorithms continue to improve, simulations of a full plasma experiment as large as the ITER scale with high fidelities, which usually combines multiple codes to resolve and connect phenomena across scales, is now a reality. WDMApp 454, for example, couples two advanced and highly scalable gyrokinetic codes, XGC and GENE (or GEM), where the former is a particle-in-cell code optimized for the treating the edge plasma while GENE (GEM) is a continuum code optimized for the core. In another example, Fusion Energy Reactor Models Integrator (FERMI) couples plasma (MHD) simulations with structural mechanics (first wall, blanket designs), neutronics, and fluid flows Badalassi et al. (2023). In practice, many ‘phenomenal approaches’ or reduced order models are adopted, which require isolation of important physics, simplified numerical algorithms, aiming for numerical results that can be directly compared with measurements or observations. Neural networks offer a completely new way to develop pragmatic reduced order models.
@@ III.1 Plasma models
As a branch of physics, the basic multi-physics framework for plasmas, which includes Newton’s mechanics, classical mechanics, Maxwell’s theory of electromagnetism, relativity, thermodynamics, statistical physics and kinetic theory, had been established by the 1950s for most practical purposes, as summarized in Fig. 8. Quantum mechanics, which has often been neglected in plasma physics, is growingly incorporated into the existing ‘classical plasma physics’ framework to better understand plasmas under extreme conditions or when the microscopic physics details of individual electrons, atoms or molecules are needed, such as in plasma chemistry. Many advances in plasma physics require modeling of very complex systems in laboratory or nature. Examples, as highlighted in Fig. 8, are magnetic fusion plasmas, space plasmas, laser plasmas, plasma accelerators, and laser fusion. One of the latest accomplishments happened at the National Ignition Facility (NIF), where nuclear fusion ignition was demonstrated in 2022. The ‘devils’ of plasma modeling have been hiding in the ‘computational details or high fidelities’ to this day. Whole device modeling of fusion plasmas, such as WDMApp, has recently become possible due to exascale computing 454. Advances in machine learning and neural networks in the 2010s, mostly due to advances outside plasma physics, is a growingly adopted in plasma physics to further enhance the traditional high-performance computing.
@@
In the context of dusty plasmas, one may divide plasma models into reactive and non-reactive plasma models. The main difference between a reactive and a non-reactive plasma lies in the significance of plasma chemistry Kushner (1988); Hollenstein (2000); Fridman (2008). In a non-reactive dusty plasma, plasma chemistry may be ignored, and dust mass, geometry, and chemical composition stay constant Pustylnik et al. (2021). We may call such models as ‘constant dust’ models. Modeling plasma chemistry can capture the material dynamics such as dust growth and destruction, meanwhile, it requires a large number of chemical species, radicals, chemical reaction cross sections and reaction rates. In industrial plasma settings Fridman (2008), including high-temperature plasma boundaries, modeling plasma chemistry is almost always necessary. Open source software, such as OpenFOAM 316, together with commercial software, such as VSimPlasma 435 and COMSOL 177, is available for the complicated physics and chemistry, and for integrated multi-disciplinary system modeling and optimization Martins and Ning (2021). In many laboratory low-temperature plasmas, non-reactive plasma assumption may be adopted for simplified computation of Newtonian dynamics and structure dynamics, while ignoring materials dynamics and plasma chemistry. For each kind of particle (electron, ions, atoms, dust grains), atomistic, kinetic, or fluid models may be used, depending on the number of particles (NN) involved and the collisional properties of the particle with the rest of system, as shown in Fig. 9. At a more fundamental level, the atomistic or first-principle models simulate individual particle motion and their interactions with others, such as particle-particle interactions, particle - electromagnetic field interactions, and particle-wall interactions. The computational cost for the best atomistic codes scales polynomially with NN as $O(N)$, which may not be practical when $N>10^{12}$ assuming Newtonian dynamics (also other parameters, such as time-step size, duration of the simulation, memory, and communications) on the state-of-the-art supercomputers, such as the Frontier supercomputer in the Oak Ridge National Laboratory 36.
@@
Besides computational cost, the selection of atomistic, kinetic or statistical, or fluid or continuum model for a specific type of particle also depends collisional properties of the particle. The collisional properties may be measured by the Knudsen number KnKn, the ratio of the collisional mean free path $\lambda$ to the characteristic length $a$. KnKn for ions interaction with a dust particle with a radius $r_{d}$ may be defined as $Kn_{d}=\lambda/r_{d}$, while $Kn_{p}=\lambda/L$ for the plasma as whole, with $L$ being the plasma size and $r_{d}\ll L$ is very common in laboratory and nature. In other words, while ions may be modeled as fluid when the problem is to quantitatively calculate the global plasma properties such as transport properties or velocity field, ion-dust interaction may need kinetic models if $Kn_{d}>1$.
@@
Attempts at balancing pragmatic computationally accessible model for a dusty plasma with scientific and logical rigor will necessarily run into the famous Hilbert’s sixth problem Slemrod (2015); Gorban (2018); Deng et al. (2025), which aimed at establishing a rigorous mathematical framework for all the laws of physics (which, of course, includes dusty plasma physics and computation), and axiomatizing the laws of physics like what has been done to geometry. From atomistic models to kinetic models, such as the Boltzmann equation, for example, the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) equations hierarchy reduces the comprehensive NN-particle description in phase space to the single-particle distribution as in the Boltzmann equation Harris (1971). Another reduction is from the kinetic equation to the fluid equations, as in Fig. 9, through the Chapman-Enskog method for charge-neutral gas. Equivalent approaches to the Chapman-Enskog method, often known as solving the closure problem, remain an active field of research in and outside plasma physics Kogelbauer and Karlin (2025).
@@
As examples of ‘atomistic’ or ‘NN-particle’ models, particle-in-cell (PIC) in combination with Monte Carlo collision (MCC), or ‘PIC-MCC’ method, is widely used in simulations of partially ionized gases Birdsall (1991); Vahedi and Surendra (1995); Alves et al. (2018); Schleede et al. (2018). The number of particles NN that can be simulated continues to increase from 103 in the 1970s to above 106 Birdsall and Langdon (2004); Verboncoeur (2005). In PIC methods, real particles (like electrons and ions) are represented by computational particles, also called macro-particles or super-particles, each representing a large number of real particles (typically 105–108Hara (2019). A weight factor is used to relate the density of computational particles to the physical density of the electrons or ions Hockney and Eastwood (1988). Axisymmetric cylindrical or 2D PIC simulations have been applied to complex boundary conditions that match those found in an experimental discharge chamber, such as a GEC cell, or to DC discharge tubes, such as the PK-4 experiment onboard the International Space Station Donkó (2011); Hartmann et al. (2020). The MCC technique is based on elementary collisions of electrons, ions and fast neutral atoms with the thermal background gas and is able to describe charged particle and energy transport at the kinetic level for individual particles. The PIC-MCC method provides a fully self-consistent, kinetic level description of the gas discharge along with space and time resolved properties (including electric potential, densities, fluxes, energy distributions, surface processes) across the entire discharge including the walls of the discharge. A hybrid-PIC-MCC (h-PIC-MCC) algorithm was described to study the effect of dust charge fluctuation on ion acoustic waves in a dusty plasma Changmai and Bora (2020).
@@
Molecular dynamics (MD) are often used to simulate NN-atom/molecule/ion/dust (microparticle) dynamics. The key idea behind a MD simulation is straightforward by using Newton’s laws of motion to trace particle motion Neyts and Brault (2017); Hollingsworth and Dror (2018). While PIC simulations track computational particles and their interactions with fields, MD simulations treat directly real particles and their interactions, including interatomic potentials such as the Lennard-Jones potential, Coulomb forces, or bonded force models. MD simulations do not require a spatial grid for force calculations unless coupled with external fields; forces are computed directly from the particle positions Frenkel and Smit (2023). One approach is to adopt existing MD codes for dusty plasmas, e.g. the microparticles were modeled using the classical MD package Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Schwabe and Graves (2013). Another approach is to develop customized MD codes for specific dusty plasma problems, e.g, molecular dynamics simulation of ion flows around microparticles was implemented in a code called molecular asymmetric dynamics (MAD) and its variants Piel (2017a); Piel et al. (2018a, b). In another example, Dynamic Response of Ions and Dust (DRIAD) is a MD code for ions and dust Matthews et al. (2020), where the dust and ion motion are resolved at different time steps, with $\Delta t=$ 10-4 s for dust and 10-9 s for ions.